OpenStudy (anonymous):

40 % out 52 deck of cards

OpenStudy (anonymous):

What exactly are you looking for? Are you asking what 40% of 52 is?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

If you take 40% and convert it to decimal you have 0.40, so... 52*0.40=??

Parth (parthkohli):

\[40\% \implies \dfrac{40}{100}\]\[\rm of \implies \times\]And so,\[\rm 40\% ~of~ 52 = \dfrac{40}{100} \times 52\]